OpenStudy (anonymous):

describe and correct the error in evaluating the function g(x)= -5x+3 when x=-3 g(-3)= -5(-3)+3 -3g=18 g= -6

OpenStudy (anonymous):

someone helpp

OpenStudy (swissgirl):

Ok I tried explaining this to someone the other day but I wasn't that clear in explaining it. Its hard to explain this concept online but let me try again

OpenStudy (swissgirl):

We are given the equation g(x)= -5x+3 This is the equation of the function g and g varies according to the value of x If g(x) was a constant such as g(x)=3 we would have a vertical line at y=3 but in our case our y keeps changing according to our x So when x=2 y=-5(2)+3 y=-10+3 y=-7 Now when x=-3 y=-5(-3)+3 y=15+3 y=18 So basically as you can see our y-value keeps changing according to what our x equals So g(-3) means what does g equal when our x equals -3

OpenStudy (anonymous):

wait i write all that?????

OpenStudy (swissgirl):

nooo But do you understand what g(-3) means?

OpenStudy (swissgirl):

The error is that they multiplied g by -3 instead of treating it as "what does g equal when x=-3" Then when x=-3 g(-3)=-5(-3)+3 g(-3)=15+3 g(-3)=18 So when x=-3 then g=18

OpenStudy (anonymous):

g(x) is a function, so g(-3) means to evaluate g(x) using x = -3 as @swissgirl demonstrated.